1.) Find an equation for the plane that passes through the origin (0, 0, 0) and is parallel to the plane - x + 3y - 2z = 6


2.) Find the distance between the point (-1, - 2, 0) and the plane 3x - y + 4z =-2


1
Expert's answer
2021-07-27T14:07:15-0400

Q1

For the line to be parallel to the line "-x+3y-2z=6" .Then,

the would be "-x+3y-2z=k" . To get k,

Since the line passes through the origin.

"k=0"

The line is therefore, "-x+3y-2z=0"

Q2

"(-1,-2,0)=(x_o,y_o,z_o)\\
The\ distance\ from\ the\\ equation\ Ax+By+Cz+D=0\ is\\
d=\dfrac{|Ax_o+By_o+Cz_o+D|}{\sqrt{A^2+B^2+C^2}}\\
A=3,B=-1,C=4,D=2\\

d=\dfrac{|3(-1)-1(-2)+4(0)+2|}{\sqrt{3^2+(-1)^2+4^2}}\\
d=\dfrac{1}{\sqrt{26}}unit"